OpenStudy (anonymous):

a force of 500 newtons will compress a spring 0.5 meters. What is the spring constant of the spring A.1000m/newton B.100m/newton C.10m/newton D..10m/newton

OpenStudy (anonymous):

\[F_{spring} = -k(x - x_0)\] \[500 = -k((-0.5) - 0) = 0.5k\]

OpenStudy (anonymous):

Use Hooke's law

OpenStudy (anonymous):

so its b right?

OpenStudy (isaiah.feynman):

Its A sir.